OpenStudy (anonymous):

how do i convert y=x^2+6x-1 into vertex form

OpenStudy (anonymous):

Hint \[y=x ^2 + 6 x +9 -9 -1 \]

OpenStudy (anonymous):

Can you finish it now?

OpenStudy (anonymous):

\[ y=x ^2 + 6 x +9 -9 -1\\ y=(x+3)^2 -9 -1\\ y=(x+3)^2 -10\\ \]

OpenStudy (anonymous):

Or \[ y+10=(x+3)^2 \]

OpenStudy (anonymous):

that helped so much.. thank you!

OpenStudy (anonymous):

YW